OpenStudy (anonymous):

Simplify the following expression. 2x - 4y + 3x2 + 7y - 16x.

OpenStudy (anonymous):

seems like you're missing a variable there somewhere. can you please type it out correctly one more time please?

OpenStudy (ybarrap):

Simplify the following: 2 x-4 y+3 x^2+7 y-16 x Group like terms in 2 x-4 y+3 x^2+7 y-16 x. Grouping like terms, 2 x-4 y+3 x^2+7 y-16 x = (7 y-4 y)+3 x^2+(2 x-16 x): (7 y-4 y)+3 x^2+(2 x-16 x) Combine like terms in 7 y-4 y. 7 y-4 y = 3 y: 3 y+3 x^2+(2 x-16 x) Combine like terms in 2 x-16 x. 2 x-16 x = -14 x: Answer: 3 y+3 x^2-14 x